Grade 12th passPhysical Chemistry

In a container of capacity 1 liter there are 10^23 molecules each of mass 10^-22g. If rms speed is 10^5cm/s then calculate the pressure of the gas

Apeksha Garg
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
rms speed => Vrms = root(3RT/M)
Now PV = nRT
or 
RT = PV/n
Vrms = root(3PV/nM)
        = root(3PV/W)
        =root(3*P*1/10^-22)
(10^5 * 10^-2 )^2 = 3*10^22*PP = 10^6/3*10^22 = > (1/3) * 10^16 atm